OpenStudy (anonymous):

simplify x+3/9/x+12/6

OpenStudy (luigi0210):

Is this your question?: \[\LARGE \frac{x+3}{\frac{9}{x}}+\frac{12}{6}\]

OpenStudy (anonymous):

no there all on top of echaother

OpenStudy (anonymous):

*eachother @Luigi0210

OpenStudy (luigi0210):

or: \[\LARGE \frac{\frac{x+3}{9}}{{x}}+\frac{12}{6}\]

OpenStudy (texaschic101):

or is it : x + (3/9)/x + 12/6

OpenStudy (anonymous):

no ais kind of like the first one but all of them are on top of eachother @Luigi0210

OpenStudy (luigi0210):

Equation editor people. Makes it easier.

OpenStudy (mathstudent55):

Maybe it's this: \(\huge \dfrac{ \frac{x+3}{9}}{\frac{x+12}{6}} \)

OpenStudy (anonymous):

@mathstudent55 yes!!!! your right

OpenStudy (mathstudent55):

If it is what I wrote, then it's a fraction over a fraction. That means a fraction divided by a fraction.

OpenStudy (mathstudent55):

\(\huge \dfrac{ \frac{x+3}{9}}{\frac{x+12}{6}}\) \( =\Large \dfrac{x+3}{9} \div \dfrac{x+12}{6} \) Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

To divide a fraction by a fraction, you multiply the first fraction by the reciprocal of the second fraction. That means keep the first fraction as it is, flip the second fraction, and multiply them together. \( =\Large \dfrac{x+3}{9} \times \dfrac {6} {x+12} \) \( =\Large \dfrac{6(x+3)}{9(x + 12)} \) Now all you need to do is to reduce the fraction. Both the 6 in the numerator and the 9 in the denominator are divisible by 3.

OpenStudy (anonymous):

so it would be 2(x+3)/3(x+12)

OpenStudy (mathstudent55):

\( =\Large \dfrac{2(x+3)}{3(x + 12)} \) You are correct.

OpenStudy (mathstudent55):

There is nothing else you can do to reduce the fraction since x + 3 and x + 12 have no common factor.

OpenStudy (mathstudent55):

In the future maybe you can use the equation editor or the draw tool to write problems like this one, so it's quicker to understand what you mean.

OpenStudy (anonymous):

whats that?

OpenStudy (mathstudent55):

This LaTex code: \huge \dfrac{ \frac{x+3}{9}}{\frac{x+12}{6}} produces this expression: \(\huge \dfrac{ \frac{x+3}{9}}{\frac{x+12}{6}}\) or you can use the draw tool: